# Deep Generative Model for Human Mobility Behavior

### v1

@@ Methods
To evaluate the ability of MobilityGen to replicate fundamental properties of individual location behavior, we compare it with representative microscopic mobility models spanning mechanistic, predictive, and generative approaches. These include: (i) classical mechanistic models such as the exploration and preferential return (EPR) model [Song et al., 2010] and the Container model [Alessandretti et al., 2020]; (ii) sequence prediction models, including a first-order Markov chain [Gambs et al., 2012] and a multi-head self-attention (MHSA)-based neural network [Hong et al., 2023]; and (iii) the deep generative MoveSim model [Feng et al., 2020]. The design and implementation details of each model are described in the following sections.
@@
    Container model. Alessandretti et al. [2020] introduced the Container model to describe how individuals allocate time and move across nested spatial scales. The model is built on a hierarchical structure $H$ with $L$ levels, where each level groups locations into increasingly broader spatial containers. Each container is associated with an attractiveness score $a$, and mobility decisions are governed by an $L\times L$ displacement matrix $D$. The probability of moving from the current location $j$ to a destination $k$ is given by:
    $p_{D_{j,h},D_{j,k}}\prod_{l\leq D_{j,k}}a(k_{l})$
@@
We evaluate the performance of all models using established location-based metrics. In addition to those reported in the main text, we include the jump length distribution [Brockmann et al., 2006] and the uncorrelated entropy [Song et al., 2010]. Results are shown in Supplementary Figure S1. To assess model fit, we compute the log-likelihood ratio, comparing the likelihood of the observed data under each model relative to MobilityGen. We additionally calculate the Wasserstein distance to quantify discrepancies between empirical and model-generated distributions for each metric (Extended Data Table 1 and Supplementary Table S1). Overall, MobilityGen provides the closest approximation to the empirical data across all evaluated metrics, except for the jump length distribution, where it tends to overestimate the frequency of longer displacements (Supplementary Figure S1d). Moreover, the median evolution of the radius of gyration across individuals, $r_{u}(k)$, over displacement steps $k$, is well described by logarithmic growth (Supplementary Figure S1b). The fit $r_{u}(k)\propto\alpha\log(k)$, with parameters $\alpha$ derived from MobilityGen ($\alpha_{\text{MobilityGen}}=2967\pm 29$), aligns most closely with the real data ($\alpha_{\text{true}}=2875\pm 35$), while the baselines exhibit notably different values ($\alpha_{\text{epr}}=6431\pm 185$, $\alpha_{\text{container}}=3784\pm 68$, $\alpha_{\text{markov}}=796\pm 26$, $\alpha_{\text{mhsa}}=1541\pm 12$, $\alpha_{\text{movSim}}=14976\pm 462$). Collectively, the results show that MobilityGen more accurately captures key dimensions of location behavior, including visitation frequency, travel distance, and spatio-temporal spread, than traditional microscopic models.
@@
To complement classical mobility metrics, we evaluate the simulated location sequences using metrics adapted from language modeling, which quantify both their similarity to and diversity from the real sequences. Similarity is measured using the Bilingual Evaluation Understudy (BLEU) score [Papineni et al., 2002], an n-gram-based metric that compares each generated sequence to its corresponding ground truth, with higher values indicating greater similarity. We implement the smoothed sentence-level version of BLEU at n-gram levels 1 through 4 (BLEU-1 to BLEU-4), commonly used in sequence generation tasks. To evaluate intra-sequence diversity, we use the distinct unigram (dist-1) metric [Gong et al., 2023], which measures the proportion of unique locations within a sequence. Lower dist-1 values indicate higher repetition and less diversity. At the sequence level, we compute the diverse 4-gram (div-4) metric [Deshpande et al., 2019], which captures the proportion of distinct 4-grams in a sequence. Higher div-4 values indicate richer generation diversity. Evaluation results for all implemented microscopic mobility models are presented in Supplementary Table S2. Overall, MobilityGen produces location visit sequences that closely mirror real-world patterns, while avoiding input replication and preserving a realistic level of behavioral diversity.
@@
While these metrics are informative, they should be interpreted with caution for the following reasons: (i) In contrast to language modeling, where multiple semantically valid reference sentences may exist, mobility datasets typically provide only one observed realization of an individual’s behavior, despite the existence of many plausible alternatives. (ii) Realistic mobility behavior involves a balance between exploration and habitual returns. As such, high diversity is not inherently desirable, and diversity metrics alone may not fully capture the plausibility of generated sequences.
Supplementary Note 2 Generating realistic activity event sequences
  • 1.
  • 2.
We assess the performance of these models using mobility metrics that characterize activity-travel patterns. In addition to the metrics reported in the main text, we introduce two supplementary measures: the number of unique location visits per day and the distribution of activity start times. Evaluation results are presented in Supplementary Figure S2. To quantify model performance, we compute the Wasserstein distance between the distributions of each metric derived from the generated activity events and those observed in the real data (Extended Data Table 3 and Supplementary Table S3). Metrics calculated from the source sequence are included as a strong reference baseline. However, it is important to account for variability in human activity behavior, as source sequences differ in both the duration and timing of activity events.
Supplementary Note 2.2 Analysis of learned attribute embeddings
We examine the structure of the learned attribute embeddings to uncover how MobilityGen internalizes and organizes behavioral attributes. Each 128-dimensional embedding is projected onto a two-dimensional space using the densMAP method, as implemented in the Python library UMAP [McInnes et al., 2018]. We use the default parameters (min_dist=0.1, dens_lambda=2.0) and confirm that alternative parameter combinations yield qualitatively consistent embedding structures. Two-dimensional visualizations are generated for both single attributes and composite activity events, computed according to Eq. 9.
Visualizations for travel mode, time, duration, and location are shown in Supplementary Figure S3. Travel modes with similar characteristics are positioned in close proximity within the embedding space; for example, walk and bike appear nearby, as do train, bus, and tram. In contrast, the “other” category appears isolated, reflecting its dissimilarity from all other modes. The time and duration embeddings in Supplementary Figure S3b and c form ordered, coherent trajectories, revealing that MobilityGen captures the continuous structure inherent in these attributes. For locations in Supplementary Figure S3d, the model organizes places according to their empirical visit frequencies in the training set, even though this frequency information is not explicitly observable from individual event sequences. This suggests that the model encodes the relative importance of locations based on usage patterns and implicitly captures the characteristic rank-frequency distribution observed in human mobility.
We next examine composite event embeddings (Supplementary Figure S4). Each point represents a mobility event defined by a combination of travel mode, time, duration, and location, with color denoting the visited location. A clear structural distinction emerges: some events form location-based clusters, while others are spatially mixed with no separation by destination. The embedding space thus loosely divides into an outer region of clustered behaviors and a central region of less location-dependent activity. This contrast suggests that the model distinguishes between different types of mobility behavior. We hypothesize that these structural differences reflect underlying activity types; for example, routine visits such as home or work (spatially consistent) versus exploratory or irregular leisure activities (spatially diffuse). Because the embedding space encodes patterns learned directly from event sequences, it provides a valuable basis for analyzing how behavioral variations relate to spatial semantics and for advancing context-aware modeling of mobility behavior.
Supplementary Note 3 Role of behavioral and contextual factors
Supplementary Note 3.1 Effect of representing activity attributes and spatial context
We conduct ablation studies on MobilityGen to assess the contributions of activity attributes and contextual information. Activity attributes serve dual roles in the model: they are embedded jointly with location data to support the learning of mobility behavior (Eq. 9), and they provide additional supervisory signals through the loss function (Eq. 18). We begin with a model trained solely on location sequences and progressively incorporate time and travel mode attributes into the pipeline. All other training configurations are held constant to ensure a fair comparison. The results are summarized in Supplementary Table S4 and S5, where model performance is measured using the Wasserstein distance between generated and real sequences across a range of mobility metrics. Each model variant is simulated three times with different random seeds, and we report the mean and standard deviation of the resulting Wasserstein distances. While the base model, trained only on historical location data, already performs strongly relative to classical mobility baselines (see performance numbers in Extended Data Table 1 and Supplementary Table S1), the inclusion of either time or travel mode data leads to further improvements. The full model, which incorporates all behavioral attributes, achieves the best overall performance, with a slight decline observed only in the jump length metric. These findings support our claim that jointly modeling behavioral attributes results in a more realistic reproduction of individual mobility patterns.
Movement-related context, including location geometry and surrounding land-use functions, is integrated with activity attributes within the embedding space. We hypothesize that this integration enhances MobilityGen’s ability to characterize locations and is particularly useful for identifying novel locations for exploration. To evaluate this, we compare the full model with a variant that excludes contextual information. As described in the main text, generated locations are grouped into three sets based on their presence in the observed sequences: Set A includes locations common to both observed and generated data, while Sets B and C consist of locations unique to the observed and generated data, respectively. We then compute the Wasserstein distance to quantify discrepancies in the distribution of visits to functional areas. We expect the model with contextual information to generate more similar visitation patterns for novel locations (i.e., comparisons between Sets B and C), and more distinct patterns between shared and novel locations (i.e., comparisons between Sets A and C), than the model without contextual information. This expectation is supported by the Wasserstein distance results in Supplementary Table S6, although the differences between models for Sets B and C are not statistically significant. This is likely due to the relatively low visit frequencies to these locations, which result in greater variability in visitation patterns.
Supplementary Note 3.2 Comparison with mobility flow prediction models
We benchmark MobilityGen against established flow models to evaluate how effectively it leverages contextual information when predicting mobility flows in unseen locations. We adopt the flow prediction task from Simini et al. [2021], where the objective is to estimate the flow $y(l_{i},l_{j})$ between each origin-destination (OD) pair $l_{i}$ and $l_{j}$ per unit time, given the total outflow $O_{i}$ from location $l_{i}$. That is, the task involves learning to distribute the total outflow from a given origin across all potential destinations based on location characteristics. Typical features include inter-location distance, population density, and contextual variables such as land use and POI types, which reflect the activity options available at each location. These correspond closely to the contextual inputs incorporated in MobilityGen.
  • 1.
    $\hat{y}(l_{i},l_{j})=O_{i}P_{ij}=O_{i}\frac{m_{j}^{\beta_{1}}f(d_{ij})}{\sum_{k}m_{k}^{\beta_{1}}f(d_{ik})}$
  • 2.
  • 3.
    Random Forest. We use the same 39 input features as Deep Gravity to train a random forest regressor for predicting OD flow intensities. Following the baseline configuration in Cabanas-Tirapu et al. [2025], we employ 1000 estimators and the default hyperparameters from scikit-learn [Pedregosa et al., 2011].
We evaluate model performance on flows originating from $\mathbb{S}_{\text{unseen}}$ using established mobility flow metrics: common part of commuters, Pearson correlation coefficient, and Jensen-Shannon divergence. Summary results are presented in Extended Data Table 4. To further assess spatial distributions, we compute the entropy of disaggregated destination probabilities. These are compared with empirical distributions and those generated by MobilityGen. Across all metrics, MobilityGen consistently achieves the highest performance and most closely reproduces the observed spatial distribution of location visits, outperforming all baseline models. We attribute these gains to the integration of detailed spatial features (e.g., urban function distributions via LDA) alongside the joint modeling of contextual and behavioral factors.
Supplementary Table S1: Supplementary performance using additional location metrics. The log-likelihood ratio measures the likelihood that the real data was generated by each model relative to MobilityGen, with positive values favoring MobilityGen (all $P<10^{-3}$, except for $P(f_{\Delta r})$ in MHSA, where $P=0.66$). The Wasserstein distance assesses the dissimilarity between generated and real location metric distributions, with higher values indicating greater divergence. Results are presented for jump length ($f_{\Delta r}$), radius of gyration ($f_{r_{u}}$), and uncorrelated entropy ($S_{\text{unc}}$); best-performing values are shown in bold.
Log-likelihood Wasserstein distance
$P(f_{\Delta r})$ $P(f_{r_{u}})$ $P(S_{\text{unc}})$ $f_{\Delta r}$ $f_{r_{u}}$ $S_{\text{unc}}$
EPR 166326 31364 74711 15882 12825 0.60
Container -6568 946 2741 1146 3578 0.20
Markov 2697 16530 329477 2223 10102 0.91
MHSA -115 4352 8355 1491 5137 0.35
MovSim 35384 79922 48838 8128 34096 0.92
MobilityGen (Ours) - - - 2318 575 0.08
@@
Supplementary Table S3: Additional metrics for behavioral alignment. The Wasserstein distance quantifies dissimilarities between metric distributions derived from simulated and real activity events. Results are presented for unique daily visited locations ($f_{\left|k\right|}^{\text{day}}$) and activity start times ($f_{t}$); best-performing values are shown in bold.
$f_{\left|k\right|}^{\text{day}}$ $f_{t}$
DITRAS 0.37 1.49
TimeGeo 1.78 1.48
MobilityGen (Ours) 0.22 0.52

### v2

@@ 4 Discussion and conclusion
To evaluate the ability of MobilityGen to replicate fundamental properties of individual location behavior, we compare it with representative microscopic mobility models spanning mechanistic, predictive, and generative approaches. These include: (i) classical mechanistic models such as the exploration and preferential return (EPR) model [Song et al., 2010b] and the Container model [Alessandretti et al., 2020]; (ii) sequence prediction models, including a first-order Markov chain [Gambs et al., 2012] and a multi-head self-attention (MHSA)-based neural network [Hong et al., 2023b]; and (iii) the deep generative MoveSim model [Feng et al., 2020]. The design and implementation details of each model are described in the following sections.
@@
    The Container model describes how individuals allocate time and move across nested spatial scales [Alessandretti et al., 2020]. The model is built on a hierarchical structure $H$ with $L$ levels, where each level groups locations into increasingly broader spatial containers. Each container is associated with an attractiveness score $a$, and mobility decisions are governed by an $L\times L$ displacement matrix $D$. The probability of moving from the current location $j$ to a destination $k$ is given by:
    $p_{D_{j,h},D_{j,k}}\prod_{l\leq D_{j,k}}a(k_{l})$ (20)
@@
Supplementary Table S1: Model performance in replicating location visit distributions. The log-likelihood ratio measures the likelihood that the real data was generated by each model relative to MobilityGen, with positive values favoring MobilityGen (all $P<10^{-3}$, except for $P(f_{\Delta r})$ in MHSA, where $P=0.66$). The Wasserstein distance assesses the dissimilarity between generated and real location metric distributions, with higher values indicating greater divergence. Results are shown for number of visits per location ($f_{k}$), jump length ($f_{\Delta r}$), radius of gyration ($f_{r_{u}}$), uncorrelated entropy ($S_{\text{unc}}$), and temporal entropy ($S_{\text{temp}}$); best-performing values are shown in bold.
Log-likelihood Wasserstein distance
$P(f_{k})$ $P(f_{\Delta r})$ $P(f_{r_{u}})$ $P(S_{\text{unc}})$ $P(S_{\text{temp}})$ $f_{k}$ $f_{\Delta r}$ $f_{r_{u}}$ $S_{\text{unc}}$ $S_{\text{temp}}$
EPR 114195 166326 31364 74711 30298 2.62 15882 12825 0.60 0.54
Container 12504 -6568 946 2741 3549 0.68 1146 3578 0.20 0.18
Markov 796846 2697 16530 329477 2728649 2.03 2223 10102 0.91 0.73
MHSA 66983 -115 4352 8355 9616 0.85 1491 5137 0.35 0.30
MovSim 268850 35384 79922 48838 22837 5.04 8128 34096 0.92 0.51
MobilityGen (Ours) - - - - - 0.26 2318 575 0.08 0.04
@@
We evaluate the performance of all models using established location-based metrics. In addition to those reported in the main text, we include the jump length distribution [Brockmann et al., 2006] and the uncorrelated entropy [Song et al., 2010a]. Results are shown in Figure S1. To assess model fit, we compute the log-likelihood ratio, comparing the likelihood of the observed data under each model relative to MobilityGen. We additionally calculate the Wasserstein distance to quantify discrepancies between empirical and model-generated distributions for each metric (Table S1). Overall, MobilityGen provides the closest approximation to the empirical data across all evaluated metrics, except for the jump length distribution, where it tends to overestimate the frequency of longer displacements (Figure S1d). Moreover, the median evolution of the radius of gyration across individuals, $r_{u}(k)$, over displacement steps $k$, is well described by logarithmic growth (Figure S1b). The fit $r_{u}(k)\propto\alpha\log(k)$, with parameters $\alpha$ derived from MobilityGen ($\alpha_{\text{MobilityGen}}=2967\pm 29$), aligns most closely with the real data ($\alpha_{\text{true}}=2875\pm 35$), while the baselines exhibit notably different values ($\alpha_{\text{epr}}=6431\pm 185$, $\alpha_{\text{container}}=3784\pm 68$, $\alpha_{\text{markov}}=796\pm 26$, $\alpha_{\text{mhsa}}=1541\pm 12$, $\alpha_{\text{movSim}}=14976\pm 462$).
Supplementary Figure S2: Comparison of spatial visitation patterns across models. Aggregated visit frequencies from real data (a) are compared with those generated by three baseline mobility models: EPR (b), Container (c), and MobilityGen (d). While all models recover broad national-scale trends, MobilityGen more accurately reproduces both high-density urban visitation and spatial dispersion into surrounding regions.
Supplementary Figure S3: Urban-level comparison of spatial visitation patterns across models. Real and simulated visit frequencies are compared across Bern (top row), Zurich (middle row), and Lucerne (bottom row). Consistent with national-level patterns, MobilityGen best captures fine-grained spatial visitation within dense urban cores and across peripheral areas.
Beyond matching marginal distributions, we also examine whether each model reproduces the geographic concentration and dispersion of visits at population scale. Figure S2 compares aggregated visitation intensity across Switzerland for the empirical data and each microscopic model, while Figure S3 provides the corresponding urban-scale comparisons for Bern, Zurich, and Lucerne. To summarize these maps quantitatively, we report the entropy of the visit-frequency distribution across locations (Table S2), where higher values indicate broader spatial dispersion. The baselines either concentrate visits too strongly or disperse them too widely relative to the data, whereas MobilityGen yields dispersion values that are closer to the empirical patterns at both national and city scales, while still showing a tendency toward slightly higher dispersion in several cases.
Supplementary Table S2: Spatial dispersion in visitation patterns. Entropy values are reported for the real data and for location traces generated by each microscopic mobility model. Higher entropy indicates broader spatial dispersion of visits. Values below those of the real data suggest insufficient reproduction of exploratory behavior, while values above indicate misalignment with population-level location preferences.
Study area Bern Zurich Luzern
EPR 1.20 3.89 5.04 3.50
Container 0.91 3.13 4.60 3.18
Markov 0.17 1.03 2.07 1.34
MHSA 0.50 2.79 4.05 2.62
MovSim 0.67 3.21 4.30 2.58
MobilityGen (Ours) 0.53 2.49 3.99 2.71
Data 0.33 1.89 3.37 2.11
Collectively, the results show that MobilityGen more accurately captures key dimensions of location behavior, including visitation frequency, travel distance, and spatio-temporal spread, than traditional microscopic models.
@@
To complement classical mobility metrics, we evaluate the simulated location sequences using metrics adapted from language modeling, which quantify both their similarity to and diversity from the real sequences. Similarity is measured using the Bilingual Evaluation Understudy (BLEU) score [Papineni et al., 2002], an n-gram-based metric that compares each generated sequence to its corresponding ground truth, with higher values indicating greater similarity. We implement the smoothed sentence-level version of BLEU at n-gram levels 1 through 4 (BLEU-1 to BLEU-4), commonly used in sequence generation tasks. To evaluate intra-sequence diversity, we use the distinct unigram (dist-1) metric [Gong et al., 2023], which measures the proportion of unique locations within a sequence. Lower dist-1 values indicate higher repetition and less diversity. At the sequence level, we compute the diverse 4-gram (div-4) metric [Deshpande et al., 2019], which captures the proportion of distinct 4-grams in a sequence. Higher div-4 values indicate richer generation diversity. Evaluation results for all implemented microscopic mobility models are presented in Table S3. Overall, MobilityGen produces location visit sequences that closely mirror real-world patterns, while avoiding input replication and preserving a realistic level of behavioral diversity.
@@
While these metrics are informative, they should be interpreted with caution: (i) In contrast to language modeling, where multiple semantically valid reference sentences may exist, mobility datasets typically provide only one observed realization of an individual’s behavior, despite the existence of many plausible alternatives. (ii) Realistic mobility behavior involves a balance between exploration and habitual returns. As such, high diversity is not inherently desirable, and diversity metrics alone may not fully capture the plausibility of generated sequences.
@@
Supplementary Note 2 Generating realistic activity event sequences
  • 1.
  • 2.
@@
Supplementary Table S4: Alignment of behavioral patterns and interdependencies. The Wasserstein distance quantifies dissimilarities between metric distributions derived from simulated and real activity events, with higher values indicating greater divergence. Source sequence results are included as a reference baseline. Results are reported for activity duration ($f_{d}$), activity start times ($f_{t}$), daily visited locations ($f_{k}^{\text{day}}$), unique daily visited locations ($f_{\left|k\right|}^{\text{day}}$), mobility motifs ($f_{\text{motifs}}$), and travel mode ($f_{m}$); best-performing values are shown in bold.
$f_{d}$ $f_{t}$ $f_{k}^{\text{day}}$ $f_{\left|k\right|}^{\text{day}}$ $f_{\text{motifs}}$ $f_{m}$
DITRAS 2.23 1.49 1.34 0.37 0.24 -
TimeGeo 1.96 1.48 1.59 1.78 1.73 -
MobilityGen (Ours) 0.47 0.52 0.29 0.22 0.14 0.048
Source sequence 0.58 1.11 0.66 0.29 0.27 0.043
We assess the performance of these models using mobility metrics that characterize activity-travel patterns. In addition to the metrics reported in the main text, we introduce two supplementary measures: the number of unique location visits per day and the distribution of activity start times. Evaluation results are presented in Figure S4. To quantify model performance, we compute the Wasserstein distance between the distributions of each metric derived from the generated activity events and those observed in the real data (Table S4). Metrics calculated from the source sequence are included as a strong reference baseline. However, it is important to account for variability in human activity behavior, as source sequences differ in both the duration and timing of activity events.
Supplementary Note 2.2 Analysis of learned attribute embeddings
We examine the structure of the learned attribute embeddings to uncover how MobilityGen internalizes and organizes behavioral attributes. Each 128-dimensional embedding is projected onto a two-dimensional space using the densMAP method, as implemented in the Python library UMAP [McInnes et al., 2018]. We use the default parameters (min_dist=0.1min\_dist=0.1, dens_lambda=2.0dens\_lambda=2.0) and confirm that alternative parameter combinations yield qualitatively consistent embedding structures. Two-dimensional visualizations are generated for both single attributes and composite activity events, computed according to Eq. 8.
Visualizations for travel mode, time, duration, and location are shown in Fig S5. Travel modes with similar characteristics are positioned in close proximity within the embedding space; for example, walk and bike appear nearby, as do train, bus, and tram. In contrast, the “other” category appears isolated, reflecting its dissimilarity from all other modes. The time and duration embeddings in Figure S5b and c form ordered, coherent trajectories, revealing that MobilityGen captures the continuous structure inherent in these attributes. For locations in Figure S5d, the model organizes places according to their empirical visit frequencies in the training set, even though this frequency information is not explicitly observable from individual event sequences. This suggests that the model encodes the relative importance of locations based on usage patterns and implicitly captures the characteristic rank-frequency distribution observed in human mobility.
We next examine composite event embeddings (Figure S6). Each point represents a mobility event defined by a combination of travel mode, time, duration, and location, with color denoting the visited location. A clear structural distinction emerges: some events form location-based clusters, while others are spatially mixed with no separation by destination. The embedding space thus loosely divides into an outer region of clustered behaviors and a central region of less location-dependent activity. This contrast suggests that the model distinguishes between different types of mobility behavior. We hypothesize that these structural differences reflect underlying activity types; for example, routine visits such as home or work (spatially consistent) versus exploratory or irregular leisure activities (spatially diffuse). Because the embedding space encodes patterns learned directly from event sequences, it provides a valuable basis for analyzing how behavioral variations relate to spatial semantics and for advancing context-aware modeling of mobility behavior.
Supplementary Note 3 Role of behavioral and contextual factors
Supplementary Note 3.1 Identifying urban functions from POI features
Supplementary Figure S7: Identifying urban functions using location LDA vectors and $k$-means clustering. a. The silhouette score assesses clustering quality, with higher values indicating more compact and better-separated classes. The score suggests that five classes yield the best separation. b. Clustering results visualized using 2D t-SNE, where each point represents a location, color-coded by class. c. Spatial distribution of the identified location classes, using the same color scheme as in b.
POIs reflect the socioeconomic activities in local regions and have been increasingly used to identify urban functions at a fine scale [Yao et al., 2017]. A well-established approach to functional identification involves obtaining a POI representation for each study unit, identifying cluster structures from these representations, and labeling the resulting classes based on expert knowledge [Niu and Silva, 2021]. Here, we perform $k$-means clustering on the LDA functional descriptor of each location. Using the silhouette score to assess clustering quality, we determine the optimal number of classes to be five (Figure S7a). Visualizing the five-class clustering result using 2-dimensional t-SNE reveals that the data points within each class are compact, while those between classes are well-separated (Figure S7b). To interpret the urban functions of each location group, we rely on the spatial distribution of classes (Figure S7c) and the POI enrichment factor for each class (Table S5). The latter quantifies the relative occurrence of different categories of POIs. For each location class $i$ and POI category $q$, the enrichment factor $EF_{i}^{q}$ is defined as:
$EF_{i}^{q}=\frac{N_{i}^{q}/N_{i}}{N^{q}/N}$ (21)
Supplementary Table S5: POI enrichment across functional clusters. First-level POI categories with more than 1,000 instances in the study area were considered. For each cluster, the top three POI types with enrichment factors greater than 1 are highlighted in bold.
POI category Cluster 0 Cluster 1 Cluster 2 Cluster 3 Cluster 4
Accommodation 1.28 0.90 0.41 0.92 0.97
Catering 1.88 0.71 0.44 0.43 0.82
Health 2.31 0.46 0.12 0.04 0.85
Leisure 0.48 1.22 3.26 0.28 1.08
Money 2.21 0.56 0.04 0.06 0.82
Natural 0.04 0.86 0.14 7.15 0.87
Others 1.01 0.83 0.87 1.81 1.03
Place of worships 0.56 1.68 0.22 0.32 0.85
Public 0.90 1.26 0.66 0.13 0.99
Shopping 1.59 0.83 0.37 0.31 0.92
Tourism 0.54 1.05 0.55 3.08 0.97
Traffic 1.00 1.09 0.64 0.50 1.02
Supplementary Note 3.2 Effect of representing activity attributes and spatial context
@@
We conduct ablation studies on MobilityGen to assess the contributions of activity attributes and contextual information. Activity attributes serve dual roles in the model: they are embedded jointly with location data to support the learning of mobility behavior (Eq. 8), and they provide additional supervisory signals through the loss function (Eq. 17). We begin with a model trained solely on location sequences and progressively incorporate time and travel mode attributes into the pipeline. All other training configurations are held constant to ensure a fair comparison. The results are summarized in Table S6 and S7, where model performance is measured using the Wasserstein distance between generated and real sequences across a range of mobility metrics. Each model variant is simulated three times with different random seeds, and we report the mean and standard deviation of the resulting Wasserstein distances. While the base model, trained only on historical location data, already performs strongly relative to classical mobility baselines (see performance numbers in Table S1), the inclusion of either time or travel mode data leads to further improvements. The full model, which incorporates all behavioral attributes, achieves the best overall performance, with a slight decline observed only in the jump length metric. These findings support our claim that jointly modeling behavioral attributes results in a more realistic reproduction of individual mobility patterns.
@@
Movement-related context, including location geometry and surrounding land-use functions, is integrated with activity attributes within the embedding space. We hypothesize that this integration enhances MobilityGen’s ability to characterize locations and is particularly useful for identifying novel locations for exploration. To evaluate this, we compare the full model with a variant that excludes contextual information. As described in the main text, generated locations are grouped into three sets based on their presence in the observed sequences: Set A includes locations common to both observed and generated data, while Sets B and C consist of locations unique to the observed and generated data, respectively. We then compute the Wasserstein distance to quantify discrepancies in the distribution of visits to functional areas. We expect the model with contextual information to generate more similar visitation patterns for novel locations (i.e., comparisons between Sets B and C), and more distinct patterns between shared and novel locations (i.e., comparisons between Sets A and C), than the model without contextual information. This expectation is supported by the Wasserstein distance results in Table S8, although the differences between models for Sets B and C are not statistically significant. This is likely due to the relatively low visit frequencies to these locations, which result in greater variability in visitation patterns.
@@
Supplementary Note 3.3 Comparison with mobility flow prediction models
We benchmark MobilityGen against established flow models to evaluate how effectively it leverages contextual information when predicting mobility flows in unseen locations. We adopt the flow prediction task with the objective to estimate the flow $y(l_{i},l_{j})$ between each origin-destination (OD) pair $l_{i}$ and $l_{j}$ per unit time, given the total outflow $O_{i}$ from location $l_{i}$ [Simini et al., 2021]. That is, the task involves learning to distribute the total outflow from a given origin across all potential destinations based on location characteristics. Typical features include inter-location distance, population density, and contextual variables such as land use and POI types, which reflect the activity options available at each location. These correspond closely to the contextual inputs incorporated in MobilityGen.
  • 1.
    $\hat{y}(l_{i},l_{j})=O_{i}P_{ij}=O_{i}\frac{m_{j}^{\beta_{1}}f(d_{ij})}{\sum_{k}m_{k}^{\beta_{1}}f(d_{ik})}$ (22)
  • 2.
  • 3.
    Random Forest. We use the same 39 input features as Deep Gravity to train a random forest regressor for predicting OD flow intensities. We employ 1000 estimators [Cabanas-Tirapu et al., 2025] and the default hyperparameters from scikit-learn [Pedregosa et al., 2011].
Supplementary Table S9: Reconstructing mobility flows at novel locations. Simulated flows are compared with real flows at locations not seen during training. For MobilityGen, individual sequences are aggregated into an origin-destination matrix to enable comparison with flow-based models. Evaluation metrics include the common part of commuters (CPC), Pearson correlation coefficient (Pearson r), and Jensen-Shannon divergence (JSD). Arrows indicate the preferred direction for each metric. Entropy quantifies the spatial dispersion of visits to novel locations. Best-performing values are shown in bold.
CPC $\uparrow$ Pearson r $\uparrow$ JSD $\downarrow$ Entropy
Gravity power-law 0.02 0.19 0.81 9.88
Random Forest 0.02 0.17 0.79 8.75
Deep Gravity 0.08 0.25 0.77 8.33
MobilityGen (Ours) 0.22 0.47 0.69 7.16
Data 6.92
We evaluate model performance on flows originating from $\mathbb{S}_{\text{unseen}}$ using established mobility flow metrics: common part of commuters, Pearson correlation coefficient, and Jensen-Shannon divergence. Summary results are presented in Table S9. To further assess spatial distributions, we compute the entropy of disaggregated destination probabilities. These are compared with empirical distributions and those generated by MobilityGen. Across all metrics, MobilityGen consistently achieves the highest performance and most closely reproduces the observed spatial distribution of location visits, outperforming all baseline models. We attribute these gains to the integration of detailed spatial features (e.g., urban function distributions via LDA) alongside the joint modeling of contextual and behavioral factors.
Supplementary Note 4 Deriving experienced income segregation
Supplementary Figure S8: Variation in income segregation across groups. Boxplots of individual-level income segregation scores for four income groups (low, lower-middle, upper-middle, and high), comparing observed data (teal) with sequences generated by TimeGeo (brown), DITRAS (purple), and MobilityGen (red). MobilityGen closely reproduces the empirical distributions and outperforms the baseline simulators in capturing variation in segregation levels across groups.
To measure income segregation, we divide individuals into four roughly equal-sized quartiles according to their reported monthly household income: low (<4,000<4,000 CHF), lower-middle (4,0008,0004,000-8,000 CHF), upper-middle (8,00012,0008,000-12,000 CHF), and high (>12,000>12,000 CHF). We then apply the experienced income segregation framework [Moro et al., 2021], described in detail below.
$S_{u}=\frac{2}{3}\sum_{q}\left|\tau_{uq}-\frac{1}{4}\right|$ (23)